# Twenty years of Nešetřil’s classification programme of Ramsey classes

### v3

@@ 2.1 Structural partition arrow
The partition arrow was originally defined for copies instead of embeddings. However, it turns out that in order to obtain Ramsey theorems, one typically needs structures to be linearly ordered for both versions of the partition arrow, and since both these notions coincide for rigid structures, there is a posteriori little difference in practice. (See Observation 2.3, Proposition 5.2, and the adjacent discussions for more details.)
@@ 2.3 Nešetřil–Rödl and Abramson–Harrington theorems
The Nešetřil–Rödl theorem, in its complete form, expands upon Theorems 2.2 and 2.4 by incorporating the ability to forbid cliques. Before stating this theorem, we first introduce the key notion of a Ramsey class.
@@
Theorem 2.5 has multiple known proofs. In addition to the partite construction, these include the original proof of Abramson and Harrington [7], as well as various applications of dual Ramsey results (see [242, Theorem 12.13] for the case of graphs, and see [15, Theorem 6.19] or [17, Section 7.2] for the general case). On the other hand, all published proofs of Theorem 2.6 use the partite construction. Recently, new alternative proofs of the Ramsey property for triangle-free graphs [77, 148] have been found (and generalized to finite binary languages [300, 15]). These new proofs extend to infinite structures, but this comes at a cost of additional complexity, and the partite construction thus remains the most effective tool for proving structural Ramsey results about finite structures to date.
@@ 2.4 Ordering property
Let $L$ be the language consisting of two binary symbols << and $\prec$, and let $\mathcal{K}$ be the class of all finite $L$-structures $\mathbf{A}$ where $(A,\prec)$ is a partial order, and (A,<)(A,{<}) is a linear extension of $(A,\prec)$. Then $\mathcal{K}$ is a Ramsey class.
@@
This theorem has an interesting history. It was announced by Nešetřil and Rödl in 1984 [224] as a result following from the partite construction. One year later, using a different method, Paoli, Trotter, and Walker proved a weaker form of this result [244, Lemma 15 and Theorem 16] [^1] This proof is sometimes considered faulty, since the paper states as Theorem 2 an infinite form of the product Ramsey theorem, which is known to hold only in its finite form. However, when Theorem 2 is applied to prove Lemma 15 and Theorem 16, the following remark saves the day: “To simplify the presentation of an argument we take $\underset{\sim}{Z}$ to be an infinite poset. Of course, we can actually choose $\underset{\sim}{Z}$ as $\underline{p}^{k}$ where $p$ is a sufficiently large integer.”. The basic idea of the proof, based on a combination of the product Ramsey theorem and the dual Ramsey theorem, was adapted by Fouché [100] to determine small Ramsey degrees of partial orders. This result directly implies Theorem 3.1 as shown by Sokić [266, Theorem 7(6)]. A self-contained presentation of this strategy was given by Solecki and Zhao [283], generalizing Theorem 3.1 to multiple linear extensions. A different approach, based on the Graham–Rothschild theorem alone, was found by Mašulović in 2018 [192] and later generalized to multiple partial orders and linear extensions jointly with Draganić [75]. In the same year, the original proof using partite construction (which is also presented here in Section A.2.1) was published by Nešetřil and Rödl [229]. Generalizations to multiple partial orders and multiple linear extensions (possibly satisfying simple axioms) can be also obtained from Theorem 3.1 and a product argument (see e. g. Bodirsky’s paper [36, Theorem 1.5 and Proposition 1.6] for a structural view of the product argument, and Arman and Rödl’s paper [9] for an independent presentation).
@@ 3.1 Observation from Lachlan’s seminar
Note that if the language of $\mathcal{K}$ contains no constants (nulary functions) then, by the hereditary property, the empty structure is a member of $\mathcal{K}$ and in this case the amalgamation property for $\mathbf{A}$ being the empty structure is exactly the joint embedding property.
@@ 2. $(b)$
Figure 2: An illustration of the proof of Observation $3.$ Copies of $\mathbf{A}$ from $\mathbf{B}_{1}$ are coloured black, copies of $\mathbf{A}$ from $\mathbf{B}_{2}$ are coloured white.
@@
The first appearance of a version of the proof of Observation 3.3 which we present here is in [214] where Nešetřil and Rödl prove that, in modern terms, if $\mathcal{C}$ is a monotone class of finite relational structures closed under inverse homomorphisms and disjoint unions and the class of all linearly ordered members of $\mathcal{C}$ is Ramsey then $\mathcal{C}$ is a free amalgamation class. The intimate connection between amalgamation and Fraïssé theory on one side, and structural Ramsey theory on the other side, was, however, only noticed by Nešetřil at Lachlan’s seminar in 1985. This then led to [206], where a more refined result is proved.
The joint embedding property is a natural assumption. Observation 3.3 thus shows us that when looking for Ramsey classes, we can restrict ourselves to ages of homogeneous structures.
@@ 3.2 Ramsey classes from homogeneous graphs
This theorem gives a complete list of the potential graph candidates for Ramseyness. By Observation 2.3, none of them is a Ramsey class due to the lack of rigidity. However, this problem can be solved by adding an order, just as in the statement of the Nešetřil–Rödl theorem:
@@
For a long time, the true role of orderings in the Ramsey property was not well understood. For example, Nešetřil in 1989 [206] only considered free orderings of the amalgamation classes given by Theorem 3.4. In this situation, the first case (the class of all ordered graphs) is a Ramsey class by the unrestricted Nešetřil–Rödl theorem (Theorem 2.5), and the second case (the class of all ordered graphs with no clique of size $k$) is a Ramsey class by the Nešetřil–Rödl theorem (Theorem 2.6). However, in the last case one only obtains a Ramsey class when putting $k=1$ or $n=1$ (see Observation 6.14), even though the Ramsey property for copies follows easily from a standard product argument (which was known to Nešetřil). Only in 2005 did the seminal paper of Kechris, Pestov, and Todorcevic [170] introduce the correct set-up for Ramsey expansions (see Section 5.2) which, in the case of disjoint unions of cliques, allows one to that restrict themselves to convex orderings (that is, orderings where every clique forms an interval). Then (possibly by further expanding by unary relations, see Example 2) one does indeed obtain Ramsey classes for the last case of Theorem 3.4 by a simple application of the product argument.
@@ 5.1 Nešetřil’s classification programme of Ramsey classes
In 2005, Nešetřil proposed a project to analyse known catalogues of homogeneous structures and initiated the classification programme of Ramsey classes (which we refer to as Nešetřil’s classification programme[207, 134]. The overall idea is symbolised in [207] as follows:
@@
    Ramsey classes $\longrightarrow$ amalgamation classes. By Observation 3.3, every hereditary isomorphism-closed Ramsey class of finite structures with the joint embedding property is an amalgamation class.
@@
Example 1 (Countable random graph).
@@ 5.2 Precompact expansions and the expansion property
Example 2 (Equivalences with two classes).
@@
Example 3 (Generic local order).
@@
Another construction of $\mathbf{S}(2)$ is to start from the order of the rationals (seen as a countable transitive tournament), randomly colour vertices with two colours so that both colour classes are dense and then reverse the direction of all arrows between vertices of different colours. In fact, this colouring is precisely the necessary Ramsey expansion. We thus consider the class of finite $L^{+}$-structures where $L^{+}$ consists of a binary relation $E$ (representing the directed edges) and a unary relation $R$ (representing one of the colour classes). The linear ordering of vertices is implicit, but can be defined based on the relations $E$ and $R$, putting $a<b$ if and only if either there is an edge from $a$ to $b$ and they belong to the same colour class, or there is an edge from $b$ to $a$ and they belong to different colour classes. The Ramsey property then follows exactly in the same way as in Example 2, and unlike in Example 2, this can be proved to be the optimal Ramsey expansion for $\mathbf{S}(2)$.
@@ 5.3 Kechris–Pestov–Todorčević correspondence
Theorem 5.1 (Kechris, Pestov, Todorčević [170, Theorem 4.8]).
@@
Theorem 5.1 is often formulated with the additional assumption that $\mathop{\mathrm{Age}}\nolimits(\mathbf{H})$ is rigid (i. e. no structure in $\mathop{\mathrm{Age}}\nolimits(\mathbf{H})$ has non-trivial automorphisms). This is however implied by our definition of a Ramsey class which colours embeddings, see Observation 2.3.
@@
Let $\mathbf{M}$ be a locally finite homogeneous structure whose age has the Ramsey property. Then $\mathop{\mathrm{Aut}}\nolimits(\mathbf{M})$ fixes a linear order, that is, there exists a linear order $\prec$ on the vertices of $\mathbf{M}$ such that for every $g\in\mathop{\mathrm{Aut}}\nolimits(\mathbf{M})$ and every $x,y\in M$ it holds that $x\prec y$ if and only if $g(x)\prec g(y)$.
@@
Theorem 5.3 (Melleray, Nguyen Van Thé, and Tsankov [204, Corollary 1.3]).
@@
Let us note that by [298], one can equivalently drop the comeagre orbit assumption in Theorem 5.3, and by essentially the same proof as [170, Theorem 10.7], one can equivalently drop the requirement that the expansion has the expansion property (see also [233]).
@@
Looking back, working hard as dogs was not truly necessary. An easy on-the-top construction to add unary functions into Ramsey classes in relational languages was later found [141, Theorem 4.29], [149, Section 7], and using this trick, the canonical Ramsey expansion of bowtie-free graphs can be obtained from the standard Nešetřil–Rödl theorem. Independently, Ramsey classes in languages containing unary functions were also considered by Sokić [271] who also approached the problem by adapting the partite construction.
@@ 6.1 Systematic approach: functions
Note that for languages without constants (nulary function), the empty structure is always a substructure; taking the free amalgam over the empty structure corresponds to finding joint embeddings.
Observation 6.3.
@@
We will prove this theorem in the appendix. Its consequence is, in particular, that if $\mathcal{K}$ is a free amalgamation class then $\mathcal{K}^{<}$ is Ramsey. The expansion property for free amalgamation classes was analyzed in [89].
@@ 6.2 Systematic approach: homomorphism-embeddings and completions
The proof of this theorem was a starting point for Nešetřil and first author’s generalizations towards classes described by forbidden homomorphisms ([71]) which they analysed earlier for homogeneous expansions [136, 135, 137, 140]. Before we state the main theorems, let us consider the following abstract question: Given an amalgamation class $\mathcal{K}$, how much additional knowledge is necessary to prove that it is a Ramsey class?
@@
Definition 6.2 (Tree amalgam [128, Definition 7.1]).
@@
Note that since we only amalgamate over substructures of irreducible substructures, it is possible to track the system of “controlled” copies of $\mathbf{A}$ in the tree amalgam and this system is a tree. We remark that [128, Definition 7.1] is slightly less restrictive as it has no requirements on $f_{i}[D]$. Since it is only used for structures $\mathbf{A}$ which contain a binary relation forming a complete graph in [128], this difference does not materialize in practice. In retrospect, it seems more practical to give a slightly more restrictive definition and avoid having additional requirements in statements using it, and so we decided to remedy it here.
@@
Let $\mathbf{D}$ be a substructure of $\mathbf{C}$ on at most $n$ vertices. By Theorem 6.6 it has a homomorphism-embedding to a tree amalgam of copies of $\mathbf{B}$, and since $\mathcal{K}$ is hereditary, has the amalgamation property, and consists of irreducible structures, Observation 6.8 gives us a completion of $\mathbf{D}$ in $\mathcal{K}$. Consequently, as $\mathcal{K}$ is a locally finite subclass of $\mathcal{R}$, we get $\mathbf{C}^{\prime}\in\mathcal{K}$ and a homomorphism-embedding $g\colon\mathbf{C}\to\mathbf{C}^{\prime}$. Note that whenever $f$ is an embedding $\mathbf{A}\to\mathbf{C}$ or $\mathbf{B}\to\mathbf{C}$ then $g\circ f$ is an embedding $\mathbf{A}\to\mathbf{C}^{\prime}$ resp. $\mathbf{B}\to\mathbf{C}^{\prime}$, since both $\mathbf{A}$ and $\mathbf{B}$ are irreducible. Consequently, a colouring of $\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{C}^{\prime})$ gives us a colouring of $\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{C})$, and as $\mathbf{C}\longrightarrow(\mathbf{B})^{\mathbf{A}}_{2}$, we obtain a monochromatic embedding $f\colon\mathbf{B}\to\mathbf{C}$. The embedding $g\circ f$ then witnesses that, indeed, $\mathbf{C}^{\prime}\longrightarrow(\mathbf{B})^{\mathbf{A}}_{2}$. ∎
@@
Fact 6.9 (Proposition 11.3 of [128], see also Proposition 2.6 of [141]).
@@
Proof of Fact 6.9 for relational $L$.

### v5

@@ 1 Introduction
Note
This document is the postprint version of the survey published in Computer Science Review [129], updated on October 31, 2025, with several added references. The reference list, however, is still far from being exhaustive.
@@ 2.1 Structural partition arrow
The partition arrow was originally defined for copies instead of embeddings. However, it turns out that in order to obtain Ramsey theorems, one typically needs structures to be linearly ordered for both versions of the partition arrow, and since both these notions coincide for rigid structures, there is a posteriori little difference in practice. (See Observation 2.5, Proposition 5.9, and the adjacent discussions for more details.)
@@ 2.3 Nešetřil–Rödl and Abramson–Harrington theorems
The Nešetřil–Rödl theorem, in its complete form, expands upon Theorems 2.4 and 2.6 by incorporating the ability to forbid cliques. Before stating this theorem, we first introduce the key notion of a Ramsey class.
@@
Theorem 2.7 has multiple known proofs. In addition to the partite construction, these include the original proof of Abramson and Harrington [7], as well as various applications of dual Ramsey results (see [246, Theorem 12.13] for the case of graphs, and see [14, Theorem 6.19] or [17, Section 7.2] for the general case). On the other hand, all published proofs of Theorem 2.9 use the partite construction. Recently, new alternative proofs of the Ramsey property for triangle-free graphs [81, 153] have been found (and generalized to finite binary languages [304, 14]). These new proofs extend to infinite structures, but this comes at a cost of additional complexity, and the partite construction thus remains the most effective tool for proving structural Ramsey results about finite structures to date.
@@ 2.4 Ordering property
Let $L$ be the language consisting of two binary symbols << and $\ll$, and let $\mathcal{K}$ be the class of all finite $L$-structures $\mathbf{A}$ where $(A,\ll)$ is a partial order, and (A,<)(A,{<}) is a linear extension of $(A,\ll)$. Then $\mathcal{K}$ is a Ramsey class.
@@
This theorem has an interesting history. It was announced by Nešetřil and Rödl in 1984 [228] as a result following from the partite construction. One year later, using a different method, Paoli, Trotter, and Walker proved a weaker form of this result [248, Lemma 15 and Theorem 16] [^1] This proof is sometimes considered faulty, since the paper states as Theorem 2 an infinite form of the product Ramsey theorem, which is known to hold only in its finite form. However, when Theorem 2 is applied to prove Lemma 15 and Theorem 16, the following remark saves the day: “To simplify the presentation of an argument we take $\underset{\sim}{Z}$ to be an infinite poset. Of course, we can actually choose $\underset{\sim}{Z}$ as $\underline{p}^{k}$ where $p$ is a sufficiently large integer.”. The basic idea of the proof, based on a combination of the product Ramsey theorem and the dual Ramsey theorem, was adapted by Fouché [104] to determine small Ramsey degrees of partial orders. This result directly implies Theorem 3.1 as shown by Sokić [270, Theorem 7(6)]. A self-contained presentation of this strategy was given by Solecki and Zhao [287], generalizing Theorem 3.1 to multiple linear extensions. A different approach, based on the Graham–Rothschild theorem alone, was found by Mašulović in 2018 [197] and later generalized to multiple partial orders and linear extensions jointly with Draganić [79]. In the same year, the original proof using partite construction (which is also presented here in Section A.2.1) was published by Nešetřil and Rödl [233]. Generalizations to multiple partial orders and multiple linear extensions (possibly satisfying simple axioms) can be also obtained from Theorem 3.1 and a product argument (see e. g. Bodirsky’s paper [36, Theorem 1.5 and Proposition 1.6] for a structural view of the product argument, and Arman and Rödl’s paper [9] for an independent presentation).
@@ 3.1 Observation from Lachlan’s seminar
Note that if the language of $\mathcal{K}$ contains no constants (nullary functions) then, by the hereditary property, the empty structure is a member of $\mathcal{K}$ and in this case the amalgamation property for $\mathbf{A}$ being the empty structure is exactly the joint embedding property.
@@ 2. $(b)$
Figure 2: An illustration of the proof of Observation 3.6. Copies of $\mathbf{A}$ from $\mathbf{B}_{1}$ are coloured black, copies of $\mathbf{A}$ from $\mathbf{B}_{2}$ are coloured white.
@@
The first appearance of a version of the proof of Observation 3.6 which we present here is in [218] where Nešetřil and Rödl prove that, in modern terms, if $\mathcal{C}$ is a monotone class of finite relational structures closed under inverse homomorphisms and disjoint unions and the class of all linearly ordered members of $\mathcal{C}$ is Ramsey then $\mathcal{C}$ is a free amalgamation class. The intimate connection between amalgamation and Fraïssé theory on one side, and structural Ramsey theory on the other side, was, however, only noticed by Nešetřil at Lachlan’s seminar in 1985. This then led to [210], where a more refined result is proved.
The joint embedding property is a natural assumption. Observation 3.6 thus shows us that when looking for Ramsey classes, we can restrict ourselves to ages of homogeneous structures.
@@ 3.2 Ramsey classes from homogeneous graphs
This theorem gives a complete list of the potential graph candidates for Ramseyness. By Observation 2.5, none of them is a Ramsey class due to the lack of rigidity. However, this problem can be solved by adding an order, just as in the statement of the Nešetřil–Rödl theorem:
@@
For a long time, the true role of orderings in the Ramsey property was not well understood. For example, Nešetřil in 1989 [210] only considered free orderings of the amalgamation classes given by Theorem 3.8. In this situation, the first case (the class of all ordered graphs) is a Ramsey class by the unrestricted Nešetřil–Rödl theorem (Theorem 2.7), and the second case (the class of all ordered graphs with no clique of size $k$) is a Ramsey class by the Nešetřil–Rödl theorem (Theorem 2.9). However, in the last case one only obtains a Ramsey class when putting $k=1$ or $n=1$ (see Observation 6.25), even though the Ramsey property for copies follows easily from a standard product argument (which was known to Nešetřil). Only in 2005 did the seminal paper of Kechris, Pestov, and Todorčević [175] introduce the correct set-up for Ramsey expansions (see Section 5.2) which, in the case of disjoint unions of cliques, allows one to that restrict themselves to convex orderings (that is, orderings where every clique forms an interval). Then (possibly by further expanding by unary relations, see Example 5.2) one does indeed obtain Ramsey classes for the last case of Theorem 3.8 by a simple application of the product argument.
@@ 5.1 Nešetřil’s classification programme of Ramsey classes
In 2005, Nešetřil proposed a project to analyse known catalogues of homogeneous structures and initiated the classification programme of Ramsey classes (which we refer to as Nešetřil’s classification programme) [211, 139]. The overall idea is symbolized in [211] as follows:
@@
    Ramsey classes $\longrightarrow$ amalgamation classes. By Observation 3.6, every hereditary isomorphism-closed Ramsey class of finite structures with the joint embedding property is an amalgamation class.
@@
Example 5.1 (Countable Random Graph).
@@ 5.2 Precompact expansions and the expansion property
Example 5.2 (Equivalences with Two Classes).
@@
Example 5.3 (Generic Local Order).
@@
Another construction of $\mathbf{S}(2)$ is to start from the order of the rationals (seen as a countable transitive tournament), randomly colour vertices with two colours so that both colour classes are dense and then reverse the direction of all arrows between vertices of different colours. In fact, this colouring is precisely the necessary Ramsey expansion. We thus consider the class of finite $L^{+}$-structures where $L^{+}$ consists of a binary relation $E$ (representing the directed edges) and a unary relation $R$ (representing one of the colour classes). The linear ordering of vertices is implicit, but can be defined based on the relations $E$ and $R$, putting $a<b$ if and only if either there is an edge from $a$ to $b$ and they belong to the same colour class, or there is an edge from $b$ to $a$ and they belong to different colour classes. The Ramsey property then follows exactly in the same way as in Example 5.2, and unlike in Example 5.2, this can be proved to be the optimal Ramsey expansion for $\mathbf{S}(2)$.
@@ 5.3 Kechris–Pestov–Todorčević correspondence
Theorem 5.8 (Kechris, Pestov, Todorčević [175, Theorem 4.8]).
@@
Theorem 5.8 is often formulated with the additional assumption that $\mathop{\mathrm{Age}}\nolimits(\mathbf{H})$ is rigid (i. e. no structure in $\mathop{\mathrm{Age}}\nolimits(\mathbf{H})$ has non-trivial automorphisms). This is however implied by our definition of a Ramsey class which colours embeddings, see Observation 2.5.
@@
Let $\mathbf{M}$ be a locally finite homogeneous structure whose age has the Ramsey property. Then $\mathop{\mathrm{Aut}}\nolimits(\mathbf{M})$ fixes a linear order, that is, there exists a linear order $\ll$ on the vertices of $\mathbf{M}$ such that for every $g\in\mathop{\mathrm{Aut}}\nolimits(\mathbf{M})$ and every $x,y\in M$ it holds that $x\ll y$ if and only if $g(x)\ll g(y)$.
@@
Theorem 5.10 (Melleray, Nguyen Van Thé, and Tsankov [208, Corollary 1.3]).
@@
Let us note that by [302], one can equivalently drop the comeagre orbit assumption in Theorem 5.10, and by essentially the same proof as [175, Theorem 10.7], one can equivalently drop the requirement that the expansion has the expansion property (see also [237]).
@@
Looking back, working hard as dogs was not truly necessary. An easy on-the-top construction to add unary functions into Ramsey classes in relational languages was later found [146, Theorem 4.29], [154, Section 7], and using this trick, the canonical Ramsey expansion of bowtie-free graphs can be obtained from the standard Nešetřil–Rödl theorem. Independently, Ramsey classes in languages containing unary functions were also considered by Sokić [275] who also approached the problem by adapting the partite construction.
@@ 6.1 Systematic approach: functions
Note that for languages without constants (nullary function), the empty structure is always a substructure; taking the free amalgam over the empty structure corresponds to finding joint embeddings.
Observation 6.4.
@@
We will prove this theorem in the appendix. Its consequence is, in particular, that if $\mathcal{K}$ is a free amalgamation class then $\mathcal{K}^{<}$ is Ramsey. The expansion property for free amalgamation classes was analysed in [93].
@@ 6.2 Systematic approach: homomorphism-embeddings and completions
The proof of this theorem was a starting point for Nešetřil and first author’s generalizations towards classes described by forbidden homomorphisms [71] which they analysed earlier for homogeneous expansions [141, 140, 142, 145]. Before we state the main theorems, let us consider the following abstract question: Given an amalgamation class $\mathcal{K}$, how much additional knowledge is necessary to prove that it is a Ramsey class?
@@
Definition 6.8 (Tree Amalgam [133, Definition 7.1]).
@@
Note that since we only amalgamate over substructures of irreducible substructures, it is possible to track the system of “controlled” copies of $\mathbf{A}$ in the tree amalgam and this system is a tree. We remark that [133, Definition 7.1] is slightly less restrictive as it has no requirements on $f_{i}[D]$. Since it is only used for structures $\mathbf{A}$ which contain a binary relation forming a complete graph in [133], this difference does not materialize in practice. In retrospect, it seems more practical to give a slightly more restrictive definition and avoid having additional requirements in statements using it, and so we decided to remedy it here.
@@
Let $\mathbf{D}$ be a substructure of $\mathbf{C}$ on at most $n$ vertices. By Theorem 6.11 it has a homomorphism-embedding to a tree amalgam of copies of $\mathbf{B}$, and since $\mathcal{K}$ is hereditary, has the amalgamation property, and consists of irreducible structures, Observation 6.17 gives us a completion of $\mathbf{D}$ in $\mathcal{K}$. Consequently, as $\mathcal{K}$ is a locally finite subclass of $\mathcal{R}$, we get $\mathbf{C}^{\prime}\in\mathcal{K}$ and a homomorphism-embedding $g\colon\mathbf{C}\to\mathbf{C}^{\prime}$. Note that whenever $f$ is an embedding $\mathbf{A}\to\mathbf{C}$ or $\mathbf{B}\to\mathbf{C}$ then $g\circ f$ is an embedding $\mathbf{A}\to\mathbf{C}^{\prime}$ resp. $\mathbf{B}\to\mathbf{C}^{\prime}$, since both $\mathbf{A}$ and $\mathbf{B}$ are irreducible. Consequently, a colouring of $\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{C}^{\prime})$ gives us a colouring of $\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{C})$, and as $\mathbf{C}\longrightarrow(\mathbf{B})^{\mathbf{A}}_{2}$, we obtain a monochromatic embedding $f\colon\mathbf{B}\to\mathbf{C}$. The embedding $g\circ f$ then witnesses that, indeed, $\mathbf{C}^{\prime}\longrightarrow(\mathbf{B})^{\mathbf{A}}_{2}$. ∎
@@
Fact 6.18 (Proposition 11.3 of [133], see also Proposition 2.6 of [146]).
@@
Proof of Fact 6.18 for Relational $L$.